### 11 Mar 2026 v2

# On automatic boundedness of operators between ordered and topological vector spaces

Abstract
We study topological boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered vector spaces to topological vector spaces. Levi and Lebesgue operators are investigated.
E. Y. Emelyanov1 and S. G. Gorokhova2
11 Sobolev Institute of Mathematics, Novosibirsk, Russia
22 Southern Mathematical Institute VSC RAS, Vladikavkaz, Russia
Keywords: ordered vector space, topological vector space, ordered Banach space, ordered Fréchet space, order-to-topology bounded operator
MSC2020: 46A40, 46B42, 47B02, 47B07, 47B60, 47B65
1 Introduction and preliminaries
Automatic boundedness of operators (enjoying some additional properties) plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g., [8, 9]). Another classical case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 7]. Apparently, the importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of such operators. The present paper is devoted to the question how far one can go with several other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 14, 15, 13, 16, 20, 18, 19, 22, 23, 25]).
Throughout, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$. We abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS). Symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net such that $\inf\limits_{\alpha}x_{\alpha}=0$, and $a,b]=\{x\in X:a\leq x\leq b\}$ for the order interval in an OVS $X$, where $a,b\in X$. Throughout, the symbol ${\cal L}(X,Y)$ stands for the space of operators between vector spaces $X$ and $Y$, and ${\rm L}(X,Y)$ for the space of bounded operators between NSs $X$ and $Y$, $B_{X}(a,r)=\{x\in X:d(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,d)$, and $B_{X}$ for $B_{X}(0,1)$ in NS $X$. It is proved in [15, Theorem 2.1] that order-to-norm bounded families of operators from a Banach lattice to an NS are equi-continuous. In [16, Theorem 2.1], it was extended to the OBS setting.
Let $X$ be a Banach space and $(Y,\tau)$ be a TVS. In absence of an additional structure in $X$, one can not say much about conditions providing boundedness of an operator $T:X\to Y$. Using of finite sets gives no proper subclass, because $T$ always carries finite sets into bounded ones. Whereas, the compactness returns us to the class of bounded operators, namely $T$ takes compact subsets into bounded sets iff $T$ is bounded. To see this, assume that $T$ is not bounded. Then, there exist an absorbing neighborhood $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ with $Tx_{n}\notin n^{2}U$ for all $n$. So, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ is not bounded.
The case of a Hilbert space in the domain was investigated in papers [18, 19], where it is proved in [19, Theorem 2.1] (see also [18, Lemma 1.2]) that
  1. -
[…]
Let us drop requirements of additional structures (like lattice operations, order, or inner products) on a space in the domain and consider:
Question 1.1.
Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it is bounded along normalized (unconditional) basic sequences of $X$?
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to Gowers and Maurey, who constructed a reflexive Banach space with no unconditional basic sequence [21]. For such a Banach space $X$, every linear functional on $X$ is bounded along bounded unconditional basic sequences simply because of absence of these sequences. So, take any unbounded functional $f:X\to\mathbb{R}$. Therefore, it makes sense to deal with the following:
Question 1.2.
How to describe pairs of Banach spaces $X$ and $Y$ such that every operator $T:X\to Y$ bounded along normalized unconditional basic sequences of $X$ is bounded?
Although Question 1.2 is interesting and challenging, we are not going to study it in this paper.
Question 1.3.
How to describe the class of Banach lattices such that every linear functional on a Banach lattice bounded along bounded disjoint sequences is bounded?
The example constructed by Storozhuk in [24] suggests that many Banach lattices (e.g., discrete Banach function spaces) do not belong this class.
The present paper deals with operators between OVSs and TVSs, so the order structure will be involved in the domain or/and codomain. We mainly focus on Banach spaces in the domain and, when suitable, Fréchet spaces. The case of Fréchet spaces in the domain occurs recently for compactly classified operators along disjoint bounded sequences in [19].
The present paper is organized as follows. Section 2 is devoted to order-to-topology continuous and order-to-topology bounded operators between OBSs (or OVSs) and TVSs. Several conditions providing topological boundedness of such operators are investigated (as well as the study undertaking of the Uniform Boundedness Principle for collectively qualified families of them). Section 3 is devoted to conditions which provide boundedness of Levi and Lebesgue operators. For example, Theorem 3.2 tells that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is topologically bounded, and Theorem 3.11 states that if an OBS in the domain has a closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded.
For further unexplained terminology and notation, we refer to [1, 3, 4].
2 On boundedness of some operators between ordered TVSs
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by many authors (see, for example, [5, 15, 16, 18, 19, 22, 23]). Some of their results were extended to operators from an OBS to a TVS in [14]. It is important to find further conditions on operators in ordered TVSs which provide automatic boundedness (see, e.g. [3, 4, 7, 13, 14]).
In this section, we continue the research initiated in [14, 15, 13, 16]. Propositions 2.3, 2.4, 2.5, and 2.6 supply with several conditions providing topological boundedness of order-to-topology continuous and order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators.
Then, we adopt the definition of $\tau$-Levi and $\tau$-Lebesgue operators from [5, 20, 14] to the OVS-setting in the domain and investigate $\tau$-Lebesgue operators from an OBS to an NS. Lemma 3.10 states that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has normal cone. Theorem 3.11 states that if the domain has closed generating normal cone then every $\sigma$-w-Lebesgue operator is bounded. Theorem 3.3 gives conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators.
This section is also aimed to introduce several notions and notation that will be used below.
Recall, that a net $(x_{\alpha})$ in an OVS $X$
  1. -
    order converges to $x$ (o-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
  2. -
    relative uniform converges to $x$ (ru-converges to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
  3. -
    is order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
  4. -
    is eventually order bounded if there exist $[a,b]$ in $X$ and $\alpha_{0}$ such that $x_{\alpha}\in[a,b]$ for all $\alpha\geq\alpha_{0}$.
  5. -
    is o-Cauchy (resp., ru-Cauchy)) if the double net $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}})$ o-converges (resp., ru-converges).
It is straightforward to see that o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) nets are eventually order bounded. In particular, o-convergent and ru-convergent (hence, o-Cauchy and ru-Cauchy) sequences are order bounded.
Definition 2.1.
Let $X$ and $Y$ be OVSs. An operator $T\in{\cal L}(X,Y)$ is
  1. $a)$
    order bounded if $T[a,b]$ is order bounded for every order interval $[a,b]$ in $X$ ((the set of such operators is denoted by ${\cal L}_{ob}(X,Y))$.
  2. $b)$
    order continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ ((shortly $T\in{\cal L}_{oc}(X,Y))$.
  3. $c)$
    ru-continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $Y$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ ((shortly $T\in{\cal L}_{rc}(X,Y))$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. An operator $T\in{\cal L}(X,Y)$ is
  1. $d)$
    order-to-topology continuous ((or $T\in{\cal L}_{o{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{onc}(X,Y)$ when $Y$ is an NS)).
  2. $e)$
    ru-to-topology continuous ((or $T\in{\cal L}_{r{\tau}c}(X,Y))$ if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ ((we write ${\cal L}_{rnc}(X,Y)$ when $Y$ is an NS)).
  3. $f)$
    order-to-topology bounded ((or $T\in{\cal L}_{o{\tau}b}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal L}_{onb}(X,Y)$ when $Y$ is an NS)).
Clearly, the sets ${\cal L}_{ob}(X,Y))$, ${\cal L}_{rc}(X,Y))$, ${\cal L}_{onc}(X,Y)$, etc., given by Definition 2.1 are vector spaces.
[…]
Definition 2.2.
Let $X$ and $Y$ be OVSs. A subset $\cal T$} of ${\cal L}(X,Y)$ is
  1. $a)$
    collectively order bounded ((or ${\cal T}\in\text{\bf L}_{ob}(X,Y)$)) whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
Let $X$ be an OVS and $(Y,\tau)$ be a TVS. A subset $\cal T$} of ${\cal L}(X,Y)$ is
  1. $b)$
    collectively order-to-topology continuous $(or{\cal T}\in\text{\bf L}_{o{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((when $Y$ is an NS, we write ${\cal T}\in\text{\bf L}_{onc}(X,Y))$.
  2. $c)$
    collectively ru-to-topology continuous $(or{\cal T}\in\text{\bf L}_{r{\tau}c}(X,Y)$)) if $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow{\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ ((we write ${\cal T}\in\text{\bf L}_{rnc}(X,Y)$ for an NS $Y)$.
  3. $d)$
    collectively order-to-topology bounded $(or{\cal T}\in\text{\bf L}_{o{\tau}b}(X,Y)$)) if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$ ((we write ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ for an NS $Y)$.
A set $\cal T$} of operators from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
  1. $e)$
    collectively continuous $(or{\cal T}\in\text{\bf L}_{c}(X,Y)$)) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c\tau}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$.
  2. $f)$
    collectively bounded $(or{\cal T}\in\text{\bf L}_{b}(X,Y)$)) if ${\cal T}U$ is $\tau$-bounded whenever $U$ is $\xi$-bounded.
The $\sigma$-versions of Definition 2.1 b–e) and Definition 2.2 b), c), and e) are obtained via replacing nets by sequences (with the corresponding notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)$, $\text{\bf L}_{c}^{\sigma}(X,Y)$, etc.).
In many cases, collectively order-to-topology bounded sets agree with collectively ru-to-topology continuous ones.
Proposition 2.3.
([14, Theorem 2.1]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{r{\tau}c}(X,Y)$, whenever $X$ is an OVS and $Y$ is a TVS. If, additionally, $X_{+}-X_{+}=X$ then $\text{\bf L}_{o{\tau}b}(X,Y)=\text{\bf L}_{r{\tau}c}(X,Y)$.
For the single operator, it follows that every order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous (and if, additionally, $X_{+}$ is generating, ${\cal L}_{o{\tau}b}(X,Y)={\cal L}_{r{\tau}c}(X,Y)$).
The following theorem is a topological version of [16, Theorem 2.1].
Proposition 2.4.
([14, Theorem 2.3]) We have $\text{\bf L}_{o{\tau}c}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$, whenever $X$ is an Archimedean OVS, $X_{+}-X_{+}=X$, and $Y$ a TVS.
In particular, every order-to-topology continuous operator from an Archimedean OVS with a generating cone to a TVS is order-to-topology bounded.
The following proposition is a partial extension of [15, Theorem 2.1] and [16, Theorem 2.8].
Proposition 2.5.
([14, Theorem 2.5]) We have $\text{\bf L}_{o{\tau}b}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$, whenever $X$ is an OBS with a closed generating cone, and $Y$ is a TVS.
It follows that every collectively order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [14], we collect few more consequences of Proposition 2.5 in the next proposition.
Proposition 2.6.
Let $X$ be an OBS with a closed generating cone.
  1. $i)$
    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets ([14, Corollary 2.9]).
  2. ii)ii)
    If the cone of $X$ is also normal and $Y$ is an NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ ([14, Corollary 2.12]).
  3. $iii)$
    If the cone of $X$ is also normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ ([14, Corollary 2.14]).
  4. $iv)$
    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded iff $Y_{+}$ is normal ([14, Proposition 2.15]).
Let $X$ be an OBS with a normal cone and $Y=(Y,\tau)$ be a Banach space with the weak topology.
  1. $v)$
    ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ ([14, Proposition 2.16]).
  2. $vi)$
    If the cone of $X$ is also closed and generating, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is bounded ([14, Corollary 2.18]).
The following theorem generalizes Proposition 2.5 to the setting of ordered Fréchet spaces. The idea of its proof below is motivated by the proof of [14, Theorem 2.5]. Recall that a Fréchet space $({\cal F},\tau)$ is an LCS, whose locally convex topology $\tau$ is metrizable by a complete metric.
Theorem 2.7.
We have $\text{\bf L}_{o{\tau}b}({\cal F},Y)\subseteq\text{\bf L}_{b}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a TVS.
Proof.
[…]
Corollary 2.8.
Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is topologically bounded.
Since $\text{\bf L}_{b}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$ whenever $X$ is an ordered TVS with a normal cone and $Y$ is a TVS, the following result follows from Proposition 2.3 and Theorem 2.7.
Corollary 2.9.
[…]
  1. $i)$
    $\text{\bf L}_{r{\tau}c}({\cal F},Y)=\text{\bf L}_{r{\tau}b}({\cal F},Y)=\text{\bf L}_{b}({\cal F},Y)$.
  2. ii)ii)
    ${\cal L}_{r{\tau}c}({\cal F},Y)={\cal L}_{o{\tau}b}({\cal F},Y)={\cal L}_{b}({\cal F},Y)$.
By [16, Theorem 2.4], ${\cal L}_{rc}(X,Y)={\cal L}_{ob}(X,Y)$ whenever OVSs $X$ and $Y$ both have generating cones. Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{o{\tau}b}(X,Y)$ if $Y$ is an ordered TVS with a normal cone, we obtain the following consequence of Corollary 2.9 ii)ii) providing conditions for automatic topological boundedness of ru-continuous operators.
Corollary 2.10.
[…]
3 On boundedness of Levi and Lebesgue operators between ordered TVSs
We use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 13, 20, 22] in the vector lattice-setting).
Definition 3.1.
An operator $T$ from an ordered TVS $(X,\tau)$ to an OVS $Y$ is
  1. $a)$
    quasi ($\sigma$-)Levi if $T$ takes $\tau$-bounded increasing nets ((sequences)) of $X$ to o-Cauchy ones (briefly, $T\in{\cal L}_{\text{\rm qLevi}}(X,Y)$ (resp., $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}(X,Y))$.
An operator $T$ from an OVS $X$ to an LCS $(Y,\xi)$ is
  1. $b)$
    $(\sigma$-))Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\xi}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
  2. $c)$
    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$ [5, Definition 1.1]. The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
  3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf. Definition 2.1d)d)).
We prove that, under a rather mild condition, quasi $\sigma$-Levi operators are topologically bounded. This result generalizes [13, Theorem 3.1].
Theorem 3.2.
Let $T$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ordered TVS $(Y,\xi)$ with a normal cone. Then $T$ is topologically bounded.
Proof.
[…]
Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ iff $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1]. The same result holds also in the following more general setting.
Theorem 3.3.
Let $T$ be a positive operator from an OVS $X$ to an ordered TVS $(Y,\tau)$ with a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{o{\tau}c}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{o{\tau}c}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Since obviously ${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{o{\tau}c}(X,Y)$.
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{o{\tau}c}(X,Y)$. Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{o{\tau}c}(X,Y)$. ∎
Corollary 3.4.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{onc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{onc})_{+}(X,Y)$.
By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue iff $T$ is w-Lebesgue. We extend this fact to OVS-setting as follows.
Theorem 3.5.
Let $T$ be a positive order continuous operator from an OVS $X$ to a Hausdorff LCS $(Y,\tau)$ ordered by a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{wLeb}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{wLeb}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$. Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$. By [4, Lemma 2.28], we conclude $Tx_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$. Thus, $T\in{\cal L}_{Leb}(X,Y)$. ∎
Corollary 3.6.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{wLeb})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{wLeb})_{+}(X,Y)$.
Question 3.7.
Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
Till the end of the section, we restrict ourselves to operators with a normed space in the co-domain. The following lemma is a direct consequence of Definition 3.1.
Lemma 3.8.
Let $X$ be an OVS and $Y$ be an LCS. Then
${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y),\ \ {\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y),$
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
${\cal L}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcap{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcup{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y).$
Lemma 3.9.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\rm L}(X,Y)$.
Proof.
The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\rm L}(X,Y)$ follows from [16, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
By [16, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded. Accordingly to [16, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous. Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
  1. (*)
    ${\cal L}_{ob}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Another case of automatic boundedness of operators concerns the order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [16, Corollary 2.2], we have:
  1. (**)
    ${\cal L}_{oc}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$. An inspection of [16, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [22] without specifying the class of (just linear or continuous linear) operators. Automatic boundedness of such operators was investigated in recent paper [23].
We are going to establish some automatic boundedness conditions for order-to-weak continuous operators on OVSs. More precisely, we shall prove Theorem 3.11 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded. We begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
Lemma 3.10.
Let $X$ be a normal OBS and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$.
Proof.
Let $T:X\to Y$ be $\sigma$-w-Lebesgue. If $T$ is not order-to-norm bounded then $T[0,u]$ is not bounded for some $u\in X_{+}$. Since $X$ is normal then $[0,u]$ is bounded, say $\sup\limits_{x\in[0,u]}\|x\|\leq M$. Pick a sequence $(u_{n})$ in $[0,u]$ with $\|Tu_{n}\|\geq n2^{n}$, and set $y_{n}:=\|\cdot\|$-$\sum\limits_{k=n}^{\infty}2^{-k}u_{k}$ for $n\in\mathbb{N}$. Then $y_{n}\downarrow\geq 0$. Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$. Since $0\leq y_{0}\leq 2^{1-n}u$ and $\|2^{1-n}u\|\leq M2^{1-n}\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. Thus, $y_{n}\downarrow 0$. It follows from $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$ that $(Ty_{n})$ is w-null, and hence is norm bounded, that is absurd because $\|Ty_{n+1}-Ty_{n}\|=\|T(2^{-n}u_{n})\|\geq n$ for all $n\in\mathbb{N}$. ∎
Combining Lemmas 3.8, 3.9 and 3.10, we obtain.
Theorem 3.11.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(E,Y)\subseteq{\rm L}(E,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(E,Y)\subseteq{\rm L}(E,Y)$.
Corollary 3.12.
(cf. [5, Proposition 3.2], [23, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
The next result is a straightforward consequence of Lemma 3.8 and Theorem 3.11.
Corollary 3.13.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\rm L}(X,Y)$.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1. c)]). From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1)].
Definition 3.14.
The topology in an ordered TVS $(X,\tau)$ (the norm in an ONS $X$) is
  1. $a)$
    order continuous if $x_{\alpha}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{\alpha}\|\to 0$) for every net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
  2. $b)$
    $(\sigma$-))order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
In certain cases o-convergence is equivalent to ru-convergence.
Lemma 3.15.
Let $X$ be an OBS with a closed normal cone.
  1. $i)$
    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
  2. ii)ii)
    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
Proof.
First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Suppose $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$. Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$, $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$. Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$. Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$.
ii)ii) The proof is similar. ∎
Theorem 3.16.
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then L^σ_Leb(X,Y)=L^σ_onc(X,Y) and L^σ_wLeb(X,Y)=L^σ_owc(X,Y).
  2. ii)ii)
    If $X$ has order continuous norm then L_Leb(X,Y)=L_onc(X,Y) and L_wLeb(X,Y)=L_owc(X,Y).
Proof.
We prove only $i)$ as the proof of ii)ii) is similar. In view of Lemma 3.8, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$. Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. Then $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$, by Lemma 3.15, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\rm L}(X,Y)$ by Theorem 3.11, and hence $\|Tx_{n}\|\to 0$. If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$. Then again $T\in{\rm L}(X,Y)$ by Theorem 3.11. So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$. ∎
The next corollary is a partial extension of results of [22, Lemma 1], [5, Lemma 2.1], [25, Theorem 3], and [23, Proposition 1.5] to the OVS setting.
Corollary 3.17.
Let $E$ be a Banach lattice with order continuous norm and $Y$ be an NS. Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and
${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
Finally, we give some conditions on order bounded operators to be order-to-norm continuous.
Theorem 3.18.
Let $X$ be an OBS with a closed normal cone and $Y$ be an ONS with a generating normal cone.
  1. $i)$
    If $X$ has $\sigma$-order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$.
  2. ii)ii)
    If $X$ has order continuous norm then ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{onc}(X,Y)$.
Proof.
We prove only $i)$ as the proof of ii)ii) is similar. Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.15, $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$. It follows from [16, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{(ru)}}0$. The normality of the cone $Y_{+}$ provides $\|Tx_{n}\|\to 0$. ∎