### 20 Apr 2026 v4

# On automatic continuity of operators from ordered to topological vector spaces

Abstract
We study continuity and boundedness of order-to-topology bounded and order-to-topology continuous operators from ordered to topological vector spaces. Several results on automatic continuity of operators from ordered Fréchet spaces to topological vector spaces are included. Levi and Lebesgue operators especially are investigated.
E. Y. Emelyanov1 and S. G. Gorokhova2
11 Sobolev Institute of Mathematics, Novosibirsk, Russia
22 Southern Mathematical Institute of the Vladikavkaz Scientific Center of the Russian Academy of Sciences, Vladikavkaz, Russia
Keywords: ordered vector space, topological vector space, Banach space, Fréchet space, order-to-topology bounded operator
MSC2020: 46A40, 46B42, 47B02, 47B07, 47B60, 47B65
1 Introduction and preliminaries
Automatic continuity of operators enjoying some additional properties plays significant role in analysis. For example, a nice theory exists on automatic continuity of algebra homomorphisms (see, e.g. [7, 8]). Another case is the automatic continuity of positive operators from Banach lattices to normed lattices [3, 4, 6]. Apparently, importance of boundedness of operators from a Banach space to a normed space lies in the Uniform Boundedness Principle for families of these operators. This paper is devoted to the study of the question how far one can go with other kinds of operators from a Banach space to a topological vector space. This question has attracted an attention in several recent papers (see, for example, [5, 11, 12, 13, 15, 14, 17, 18, 21, 22, 24]).
Throughout, vector spaces are real, and operators are linear. When suitable, we abbreviate a topological vector space (normed space, locally convex space, ordered vector space, ordered normed space, ordered Banach space) by TVS (NS, LCS, OVS, ONS, OBS respectively). In what follows, symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$; $a,b]=\{x\in X:a\leq x\leq b\}$ for an order interval in an OVS $X$; ${\cal L}(X,Y)$ for the space of operators between vector spaces $X$ and $Y$; $B_{X}(a,r)=\{x\in X:p(x,a)\leq r\}$ for the closed ball centered at $a\in X$ of radius $r\geq 0$ in a metric space $(X,p)$; and $B_{X}:=B_{X}(0,1)$, where $X$ is a NS. It is proved in [13, Theorem 2.1] that each order-to-norm bounded family of operators from a Banach lattice to an NS is equi-continuous. In [15, Theorem 2.1], this result is extended to the OBS setting. In [12], the further extension is obtained in the case of a TVS in the co-domain. In the present paper, we prove Theorem 2.6 that generalizes [12, Theorem 2.1] to operators from an ordered Fréchet space with a closed generating cone to a TVS.
Let $T:X\to Y$ be an operator from a normed space $X$ to a TVS $(Y,\tau)$. In absence of additional structures in $X$, one cannot say much about conditions providing topological boundedness/continuity of $T$. For example, using of finite sets gives no proper subclass, because $T$ always carries finite sets into $\tau$-bounded sets. On the other hand, usage of compactness returns us to bounded operators. Indeed, if $T$ is $\tau$-bounded then $T$ takes compact sets into bounded sets. On the other hand, if $T$ is not $\tau$-bounded then there exist a circled $U\in\tau(0)$ and a sequence $(x_{n})$ in $B_{X}$ such that $Tx_{n}\notin n^{2}U$ for all $n\in\mathbb{N}$. In this case, the image $\big\{T(\frac{x_{n}}{n})\big\}_{n=1}^{\infty}\cup\{0\}$ of the compact set $\big\{\frac{x_{n}}{n}\big\}_{n=1}^{\infty}\cup\{0\}$ under $T$ is not $\tau$-bounded.
The case of a Hilbert space in the domain was investigated in papers [17, 18], where it is proved in [18, Theorem 2.1] (see also [17, Lemma 1.2]) that:
  1. -
[…]
Let us drop requirements of additional structures (e.g., lattice operations, order, or inner products) on a space in the domain and consider:
Question 1.1.
Is an operator $T:X\to Y$ between Banach spaces $X$ and $Y$ bounded, whenever it takes normalized (unconditional) basic sequences of $X$ onto bounded sequences of $Y$?
The ‘‘conditional part’’ of Question 1.1 is open even when $X$ is reflexive, whereas the ‘‘unconditional part’’ has a negative answer due to the example of a reflexive Banach space with no unconditional basic sequence constructed by Gowers and Maurey [20] (every linear functional on this space is bounded on normalized unconditional basic sequences simply because of absence of such sequences, so, we take any unbounded functional on this space). Therefore, it makes sense to deal with the following questions.
  1. -
    Under which conditions on Banach spaces $X$ and $Y$ every operator $T:X\to Y$ that is bounded on normalized unconditional basic sequences of $X$ is bounded?
  2. -
    How to describe the class $S$ of Banach lattices such that every linear functional on a Banach lattice $X$ from $S$ is bounded, whenever it is bounded on every normalized disjoint sequence of $X$? An example, recently constructed by Storozhuk in [23] suggests that many of Banach lattices (e.g., infinite dimensional discrete Banach function spaces) do not belong to the class $S$.
The present paper deals with operators between vector spaces, where the order structure is assumed in the domain or/and codomain. We focus on Banach spaces and, when suitable, Fréchet spaces in the domain. The case of Fréchet spaces in the domain occurred recently in study of compactly classified operators along disjoint bounded sequences [18].
This paper is organized as follows. In Section 2, we survey briefly some recently published results on order-to-topology continuous/bounded operators between OBSs and TVSs. Then, we prove Theorem 2.6 on continuity of order-to-topology bounded operators from an ordered Fréchet space with a closed generating cone to a TVS. In Section 3, we study conditions providing continuity of Levi and Lebesgue operators. For instance, we prove: Theorem 3.2 telling us that each quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered TVS with a normal cone is continuous; and Theorem 3.12 saying that each $\sigma$-w-Lebesgue operator from an ordered Banach space with a closed generating normal cone to a normed space is bounded.
For further unexplained terminology and notation, we refer the reader to [1, 3, 4].
2 On continuity of some operators between ordered topological vector spaces
Order-to-topology continuous and order-to-norm bounded operators have been studied recently (mostly in the Banach lattice setting) by several authors (see, e.g., [5, 13, 15, 17, 18, 21, 22]). Some of their results were extended to operators from an OBS to a TVS in [12]. It is important to find further conditions on operators in ordered TVSs which provide automatic continuity/boundedness. In this section, we continue the research initiated in [12, 13, 11, 15]. First, we point out Propositions 2.4 and 2.5 providing several conditions for topological boundedness of order-to-topology continuous and of order-to-norm bounded operators, and the Uniform Boundedness Principle for collectively qualified families of such operators. Then, we prove Theorem 2.6 telling us that every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is continuous. We need several further notions and notation.
Definition 2.1.
A net $(x_{\alpha})$ in an ordered vector space $X$ is
  1. $a)$
    order convergent to $x$ (o-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{o}}x$) if there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there is $\alpha_{\beta}$ such that $\pm(x_{\alpha}-x)\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$.
  2. $b)$
    relative uniform convergent to $x$ (ru-convergent to $x$, or $x_{\alpha}{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $\pm(x_{\alpha}-x)\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$.
  3. $c)$
    order bounded if there exists an order interval $[a,b]$ in $X$ such that $x_{\alpha}\in[a,b]$ for all $\alpha$.
  4. $d)$
    eventually order bounded if there exist $a,b\in X$ and $\alpha_{0}$ with $x_{\alpha}\in[a,b]$ for $\alpha\geq\alpha_{0}$.
  5. $e)$
    o-Cauchy (resp., ru-Cauchy) if $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{o}}0$ (resp., $(x_{\alpha^{\prime}}-x_{\alpha^{\prime\prime}}){\xrightarrow[]{ru}}0$).
It is straightforward to see that every o-convergent and every ru-convergent (hence, every o-Cauchy and every ru-Cauchy) net is eventually order bounded. In particular, o-convergent and ru-convergent (hence, also o-Cauchy and ru-Cauchy) sequences are order bounded.
Definition 2.2.
An operator $T$ from an OVS $X$ to an OVS $Y$ is
  1. $a)$
    order bounded $(T\in{\cal L}_{ob}(X,Y))$ if $T[a,b]$ is order bounded for every $a,b]\subseteq X$.
  2. $b)$
    order continuous $(T\in{\cal L}_{oc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{o}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{o}}0$.
  3. $c)$
    ru-continuous $(T\in{\cal L}_{rc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{ru}}0$ in $Y$, whenever $x_{\alpha}{\xrightarrow[]{ru}}0$.
An operator $T$ from an OVS $X$ to a TVS $(Y,\tau)$ is
  1. $d)$
    order-to-topology continuous $(T\in{\cal L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
  2. $e)$
    ru-to-topology continuous $(T\in{\cal L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow Tx_{\alpha}{\xrightarrow[]{\tau}}0$.
  3. $f)$
    order-to-topology bounded $(T\in{\cal L}_{otb}(X,Y))$ if $T[a,b]$ is $\tau$-bounded for every $a,b\in X$.
An operator $T$ from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
  1. $g)$
    continuous $(T\in{\cal L}_{tc}(X,Y))$ if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
  2. $e)$
    topologically-bounded $(T\in{\cal L}_{b}(X,Y))$ if $TU$ is $\tau$-bounded for every $\xi$-bounded $U$.
Evidently, the sets ${\cal L}_{ob}(X,Y)$, ${\cal L}_{otc}(X,Y)$, etc. given by Definition 2.2 are vector spaces.
We recall several notions on collectively qualified sets of operators (cf. [9, 10, 12, 13, 17, 18, 15]). A family ${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ of nets indexed by a directed set $A$ in a TVS $(X,\tau)$ collective $\tau$-converges to 0 (shortly, ${\cal B}{\xrightarrow[]{c\tau}}0$) if, for each $U\in\tau(0)$, there exists $\alpha_{U}$ such that $x^{b}_{\alpha}\in U$ for all $\alpha\geq\alpha_{U}$ and $b\in B$.
Definition 2.3.
A set $\cal T$} of operators from an OVS $X$ to an OVS $Y$ is
  1. $a)$
    equi-order bounded $({\cal T}\in\text{\bf L}_{ob}(X,Y))$, whenever ${\cal T}[a,b]$ is order bounded for every $a,b\in X$.
A set $\cal T$} of operators from an OVS $X$ to a TVS $(Y,\tau)$ is
  1. $b)$
    equi-order-to-topology continuous $({\cal T}\in\text{\bf L}_{otc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{o}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
  2. $c)$
    equi-ru-to-topology continuous $({\cal T}\in\text{\bf L}_{rtc}(X,Y))$ if $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow{\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$.
  3. $d)$
    equi-order-to-topology bounded $({\cal T}\in\text{\bf L}_{otb}(X,Y))$ if ${\cal T}[a,b]$ is $\tau$-bounded for every $a,b\in X$.
A set $\cal T$} of operators from a TVS $(X,\xi)$ to a TVS $(Y,\tau)$ is
  1. $e)$
    equi-continuous $({\cal T}\in\text{\bf L}_{tc}(X,Y))$ if ${\cal T}x_{\alpha}{\xrightarrow[]{c\tau}}0$, whenever $x_{\alpha}{\xrightarrow[]{\xi}}0$.
  2. $f)$
    equi-topologically-bounded $({\cal T}\in\text{\bf L}_{b}(X,Y))$ if ${\cal T}U$ is $\tau$-bounded, whenever $U$ is $\xi$-bounded.
The sequential versions of Definition 2.2 b)–e), g) and Definition 2.3 b), c), and e) are obtained via replacing nets by sequences (with notations ${\cal L}_{oc}^{\sigma}(X,Y)$, ${\cal L}_{otc}^{\sigma}(X,Y)$, $\text{\bf L}_{tc}^{\sigma}(X,Y)$, etc.).
We recall several recent results concerned the notions listed in Definitions 2.2 and 2.3.
Proposition 2.4.
([12, Theorem 2.1]) If $X$ is an OVS and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{rtc}(X,Y)$. If additionally $X_{+}$ is generating then $\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{rtc}(X,Y)$.
In particular, each order-to-topology bounded operator from an OVS $X$ to a TVS $Y$ is ru-to-topology continuous, and if additionally $X_{+}$ is generating, then ${\cal L}_{otb}(X,Y)={\cal L}_{rtc}(X,Y)$. The following proposition is a partial extension of [13, Theorem 2.1] and [15, Theorem 2.8].
Proposition 2.5.
([12, Theorem 2.5]) If $X$ is an OBS with a closed generating cone and $Y$ is a TVS then $\text{\bf L}_{otb}(X,Y)\subseteq\text{\bf L}_{b}(X,Y)$.
It follows that every equi-order-to-norm bounded set of operators from an OBS with a closed generating cone to an NS is uniformly bounded. In particular, each order bounded operator from an OBS with a closed generating cone to an ONS with a normal cone is continuous. Following [12], we collect few more statements below.
Let $X$ be an ordered Banach space with a closed generating cone.
  1. $i)$
    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets [12, Corollary 2.9].
  2. ii)ii)
    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rtc}(X,Y)=\text{\bf L}_{otb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
  3. $iii)$
    If $X_{+}$ is normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ [12, Corollary 2.14].
  4. $iv)$
    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, where $X^{\sim}$ is an order dual of $X$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded if and only if $Y_{+}$ is normal [12, Proposition 2.15].
  5. $v)$
    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ [12, Proposition 2.16].
  6. $vi)$
    If $X_{+}$ is closed generating and normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is topologically bounded [12, Corollary 2.18].
[…]
Theorem 2.6.
$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{tc}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
Proof.
Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{tc}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
$V_{n}:=B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}-B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}.$
Clearly, $\{B_{\cal F}(0,2^{-2n}):n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of $\tau$-closed sets and $B_{\cal F}(0,2^{-2(n+1)})+B_{\cal F}(0,2^{-2(n+1)})\subseteq B_{\cal F}(0,2^{-2n})$ holds for each $n$. Furthermore, each $B_{\cal F}(0,2^{-2n})$ is circled. To see this, it suffices to show that $p(x,0)\leq 2^{-2n}$ implies $p(tx,0)\leq 2^{-2n}$ whenever $|t|\leq 1$. So, let $p(x,0)\leq 2^{-2n}$ and $|t|\leq 1$. Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$. By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of circled sets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
[…]
Corollary 2.7.
Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a topological vector space is continuous.
Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ if $Y$ is a TVS with a normal cone [4, Lemma 2.22], we obtain the following conditions for automatic continuity of order bounded operators.
Corollary 2.8.
Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
3 On continuity of the Levi and the Lebesgue operators
In this section, we study $\tau$-Levi and $\tau$-Lebesgue operators [5, 19, 12] in the more general OVS-setting. We prove Theorem 3.2 saying that a quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous. This theorem is a far going generalization of [11, Theorem 3.1], where the continuity is established for a quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice. We mention also: Lemma 3.11 saying that each $\sigma$-w-Lebesgue operator is order-to-norm bounded whenever an OBS in the domain has a normal cone; Theorem 3.12 saying that, if the domain has closed generating normal cone, then every $\sigma$-w-Lebesgue operator is bounded; and Theorem 3.3 providing conditions for ($\sigma$-)order to (weak) norm continuity of ($\sigma$-)(w-)Lebesgue operators. We shall use below the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 11, 19, 21] in the vector lattice-setting).
Definition 3.1.
An operator $T$ from an ordered TVS $(X,\tau)$ to an OVS $Y$ is
  1. $a)$
    quasi $\sigma$-Levi if $T$ takes $\tau$-bounded increasing sequences of $X$ to o-Cauchy ones.
An operator $T$ from an OVS $X$ to an LCS $(Y,\xi)$ is
  1. $b)$
    $(\sigma$-))Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\tau}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$. The set of such operators is denoted by ${\cal L}_{Leb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{Leb}(X,Y)$).
  2. $c)$
    $(\sigma$-))w-Lebesgue if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\downarrow 0$ in $X$. The set of such operators is denoted by ${\cal L}_{wLeb}(X,Y)$ (resp., ${\cal L}^{\sigma}_{wLeb}(X,Y)$).
  3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.2d)d)).
First, we prove that under rather mild conditions, a quasi $\sigma$-Levi operator is continuous. This generalizes [11, Theorem 3.1] which states that every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
Theorem 3.2.
Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered topological vector space with a normal cone is continuous.
Proof.
Let $T:{\cal F}\to Y$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ONS $Y$ with a normal cone. Toward a contradiction, suppose that $T$ is not continuous.
[…]
Letting $x_{n}:=2^{n}z_{n}$, denote $u_{i}=\sum\limits_{k=1}^{i}x_{k}$ for each $i\in\mathbb{N}$. Clearly, $(u_{i})$ is an increasing sequence. Take an $l\in\mathbb{N}$. It is straightforward to see that, for each $m\in\mathbb{N}$,
$\sum\limits_{k=l+1}^{m+l+1}x_{k}=\sum\limits_{k=l+1}^{m+l+1}2^{k}z_{k}\in\sum\limits_{k=l+1}^{m+l+1}2^{k}B_{\cal F}(0,2^{-2k})\subseteq\sum\limits_{k=0}^{m}2^{-k}B_{\cal F}(0,2^{-(l+1)})\subseteq B_{\cal F}(0,2^{-l}).$
Find an $N>0$ satisfying $\sum\limits_{k=1}^{j}x_{k}\in NB_{\cal F}(0,2^{-l})$ for j=1,,lj=1,...,l. Then, for every $i\in\mathbb{N}$,
$u_{i}=\sum\limits_{k=1}^{i}x_{k}=\sum\limits_{k=1}^{\min(i,l)}x_{k}+\sum\limits_{k=\min(i,l)+1}^{i}x_{k}\in(N+1)B_{\cal F}(0,2^{-l}).$
As $\{B_{\cal F}(0,2^{-l})\}_{l\in\mathbb{N}}$ is a $\tau$-neighborhood base of zero, the sequence $(u_{i})$ is $\tau$-bounded.
As $T$ is a quasi $\sigma$-Levi operator, the sequence $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded. Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is topologically bounded by [4, Lemma 2.22]. It is absurd because $Tu_{i+1}-Tu_{i}=Tx_{i+1}=2^{i+1}Tz_{i+1}\notin 2^{i+1}W$ for all $i\in\mathbb{N}$. ∎
Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ if and only if $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1]. The same result holds also in the following more general setting.
Theorem 3.3.
Let $T$ be a positive operator from an ordered vector space $X$ to an ordered topological vector space $(Y,\tau)$ with a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{otc}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{otc}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Since obviously ${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{otc}(X,Y)$.
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$. Then, there exist a net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{otc}(X,Y)$. ∎
Corollary 3.4.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{otc})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{otc})_{+}(X,Y)$.
By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue if and only if $T$ is w-Lebesgue. We extend this fact to the OVS-setting as follows.
Theorem 3.5.
Let $T$ be a positive order continuous operator from an OVS $X$ to a Hausdorff LCS $(Y,\tau)$ ordered by a normal cone. Then
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{wLeb}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{wLeb}(X,Y).$
Proof.
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Suppose $T\in{\cal L}_{wLeb}(X,Y)$ and take $x_{\alpha}\downarrow 0$ in $X$. Since $T\geq 0$, we have $Tx_{\alpha}\downarrow\geq 0$ and, as $T$ is w-Lebesgue, $Tx_{\alpha}{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$. By [4, Lemma 2.28], we conclude $Tx_{\alpha}{\xrightarrow[]{\tau}}0$. Thus, $T\in{\cal L}_{Leb}(X,Y)$. ∎
Corollary 3.6.
Let $X$ be an OVS and $Y$ be a normal ONS. Then $({\cal L}_{Leb})_{+}(X,Y)=({\cal L}_{wLeb})_{+}(X,Y)$ and $({\cal L}^{\sigma}_{Leb})_{+}(X,Y)=({\cal L}^{\sigma}_{wLeb})_{+}(X,Y)$.
Question 3.7.
Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
The following lemma is a direct consequence of Definition 3.1.
Lemma 3.8.
Let $X$ be an OVS and $Y$ be an LCS. Then
${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y),\ \ {\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y),$
${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
${\cal L}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcap{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{wLeb}(X,Y)\bigcup{\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}^{\sigma}_{wLeb}(X,Y).$
Lemma 3.9.
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then, ${\cal L}_{otb}(X,Y)={\cal L}_{b}(X,Y)$.
Proof.
The inclusion ${\cal L}_{otb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
Another case of automatic continuity of operators concerns their order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have: ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal. Condition $X_{+}-X_{+}=X$ is essential here since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$.
Example 3.10.
([15, Example 2.12 b)]) Choose a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of an infinite dimensional Banach space $X$ and identify $X$ with $c_{00}(\Xi)$. Consider a positive cone $X_{+}:=\Big\{\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big\}$ in $X$. The cone $X_{+}$ is generating. However $X_{+}$ is not closed ((e.g., by [6, Proposition 4.1])). Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$. It is straightforward to see that the operator $T$ is order-to-norm bounded and positive, but $T$ is not bounded.
It is worth nothing that the order-to-norm and order-to-weak continuous operators on vector lattices were defined in [21] without specifying the class of (just linear or continuous linear) operators. Automatic boundedness of such operators was investigated in the recent paper [22].
We are going to establish further conditions for automatic boundedness of order-to-weak continuous operators between ordered Banach spaces and and normed spaces. More precisely, we prove Theorem 3.12 which tells us that every $\sigma$-w-Lebesgue operator from an OBS with a closed generating normal cone to an NS is bounded. Let us begin with conditions under which $\sigma$-w-Lebesgue operators are order-to-norm bounded.
Lemma 3.11.
Let $X$ be an ordered Banach space with a normal cone and $Y$ a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$. In particular, ${\cal L}^{\sigma}_{owc}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$.
Proof.
Let $T:X\to Y$ be $\sigma$-w-Lebesgue. If $T$ is not order-to-norm bounded then $T[0,u]$ is not bounded for some $u\in X_{+}$. Since $X$ is normal then $[0,u]$ is bounded, say $\sup\limits_{x\in[0,u]}\|x\|\leq M$. Pick a sequence $(u_{n})$ in $[0,u]$ with $\|Tu_{n}\|\geq n2^{n}$, and set $y_{n}:=\|\cdot\|$-$\sum\limits_{k=n}^{\infty}2^{-k}u_{k}$ for $n\in\mathbb{N}$. Then $y_{n}\downarrow\geq 0$. Let $0\leq y_{0}\leq y_{n}$ for all $n\in\mathbb{N}$. Since $0\leq y_{0}\leq 2^{1-n}u$ it follows that $\|y_{0}\|\leq M2^{1-n}$ for every $n$. Thus, $y_{n}\downarrow 0$. It follows from $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$ that $(Ty_{n})$ is w-null, and hence is norm bounded, that is absurd because $\|Ty_{n+1}-Ty_{n}\|=\|T(2^{-n}u_{n})\|\geq n$ for all $n\in\mathbb{N}$. ∎
Combining Lemmas 3.8, 3.9, and 3.11, we obtain.
Theorem 3.12.
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
Corollary 3.13.
(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an normed space is bounded.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded [5, Remark 1.1. c)]. From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].
Definition 3.14.
The norm in an ordered normed space $X$ is order continuous if $\|x_{\alpha}\|\to 0$ for every net $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
In certain cases o-convergence is equivalent to ru-convergence.
Lemma 3.15.
Let $X$ be an OBS with a closed normal cone and the norm in $X$ is order continuous. Then, $x_{\alpha}{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}{\xrightarrow[]{ru}}0$ in $X$.
Proof.
Since $X_{+}$ is normal, $X$ is Archimedean, and hence $x_{\alpha}{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}{\xrightarrow[]{o}}0$ in $X$.
Let $X$ has order continuous norm and $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Since the norm in $X$ is order continuous and $X_{+}$ is normal, there exists an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$. Set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}{\xrightarrow[]{ru}}0$. ∎
Theorem 3.16.
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. If $X$ has order continuous norm then
${\cal L}_{Leb}(X,Y)={\cal L}_{otc}(X,Y)\ \ and\ \ {\cal L}_{wLeb}(X,Y)={\cal L}_{owc}(X,Y).$
Proof.
In view of Lemma 3.8, we need to prove ${\cal L}_{Leb}(X,Y)\subseteq{\cal L}_{otc}(X,Y)$ and ${\cal L}_{wLeb}(X,Y)\subseteq{\cal L}_{owc}(X,Y)$. Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. Then $x_{\alpha}{\xrightarrow[]{ru}}0$ by Lemma 3.15, and hence $\|x_{\alpha}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}_{Leb}(X,Y)$ then $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$. By Theorem 3.12, $T\in{\cal L}_{b}(X,Y)$, and hence $\|Tx_{\alpha}\|\to 0$. Thus, $T\in{\cal L}_{otc}(X,Y)$. If $T\in{\cal L}_{wLeb}(X,Y)$ then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12. Thus, $\|Tx_{\alpha}\|\to 0$, and hence $Tx_{\alpha}{\xrightarrow[]{w}}0$. So, $T\in{\cal L}_{owc}(X,Y)$. ∎
The next corollary is a partial extension of results of [21, Lemma 1], [5, Lemma 2.1], [24, Theorem 3], and [22, Proposition 1.5] to the OVS setting.
Corollary 3.17.
Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space. Then, ${\cal L}_{Leb}(E,Y)={\cal L}_{otc}(E,Y)$ and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
Finally, we give the conditions for order-to-topology continuity of order bounded operators.
Theorem 3.18.
Let $X$ be an ordered Banach space with a closed normal cone and order continuous norm, $Y$ be an ordered normed space with a generating normal cone, and $T:X\to Y$ and order bounded operator. Then $T$ is order-to-topology continuous.
Proof.
Let $x_{\alpha}{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.15, $x_{\alpha}{\xrightarrow[]{ru}}0$. It follows from [15, Corollary 2.5] that $Tx_{\alpha}{\xrightarrow[]{ru}}0$. The normality of the cone $Y_{+}$ ensures $\|Tx_{\alpha}\|\to 0$. ∎